OpenStudy (anonymous):

Evaluate the Integral : Integral from (1/2) to (sqrt(3)/2) of (6/(1-t^2))dt I'm in calc 2, What U-substitution should I use?

OpenStudy (amistre64):

t - sin

OpenStudy (amistre64):

t = sin

OpenStudy (amistre64):

maybe not; its hard to read the equation

OpenStudy (anonymous):

\[\int\limits_{1/2}^{\sqrt(3)/2} (6/(1-t^2))dt\]

OpenStudy (amistre64):

\[ \int_{1/2}^{\sqrt{3}/2} \frac{6}{(1-t^2)}dt\]

OpenStudy (anonymous):

Yes like that

OpenStudy (amistre64):

split the denom up into its factors and do partials

OpenStudy (anonymous):

6/(1+t) 6/(1-t) ?

OpenStudy (amistre64):

yep ... or, try it as:\[6\int \frac{1}{1-sin^2(u)}du\] \[6\int \frac{1}{cos^2(u)}du\] \[6\int sec^2(u)du\]

OpenStudy (amistre64):

im missing something arent i

OpenStudy (anonymous):

Well the denominator already looks like a formula I have already -- \[\int\limits ( du / (a^2+u^2) ) = 1/a * \arctan(u/a)+c\]

OpenStudy (amistre64):

t = sin(u) dt = cos(u) du \[6\int_{1/2}^{\sqrt{3}/2} \frac{1}{(1-t^2)}dt\] \[6\int_{1/2}^{\sqrt{3}/2} \frac{cos(u)}{(1-sin^2(u))}du\] \[6\int_{1/2}^{\sqrt{3}/2} \frac{cos(u)}{cos^2(u)}du\] \[6\int_{1/2}^{\sqrt{3}/2} \frac{1}{cos(u)}du\] \[6\int_{1/2}^{\sqrt{3}/2} sec(u)du\]

OpenStudy (anonymous):

so I just need to get rid of the 6.. oh wait, im dumb, can i pull it out by constant factor

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

\[\frac{6}{(1-t^2)}=\frac{6}{(1-t)(1+t)}\] \[\frac{6}{(1-t)(1+t)}=\frac{A}{(1-t)}+\frac{B}{1+t}\] \[6=A(1+t)+B(1-t)\] \[t=1;\ 6=A(1+1)+B(1-1)\] \[t=1;\ 6=A(2);\ A = 3\] \[t=-1;\ 6=A(1-1)+B(1+1)\] \[t=-1;\ 6=B(2);\ B=3\] \[\int \frac{3}{1-t}dt\ +\int\frac{3}{1+t}dt\] right?

OpenStudy (anonymous):

With your answer, I get kind of confused -- Usually we change the limits, and I realize its just the same thing but when you went from du to dt, you have to add a arccos, correct?

OpenStudy (amistre64):

and arccos no; i just had to adjust for the chain rule that pops out the cos(u) up top

OpenStudy (anonymous):

Okay, i get it, Because Dt = cos(u) du , they're both put in, without countering anythign on the other side.. Now how do I go about the resub, by putting in arcsin(t) in as u?

OpenStudy (amistre64):

hmm ... ive never considered resubing it at this point; since it is at its simplest to "work" with

OpenStudy (anonymous):

Well the directions are to evaluate the integral rather than simplify, if thats how you're looking at it.

OpenStudy (amistre64):

by turning it into something alot simpler to solve; you make the work of integration easier and doable; the method used "u-sub", "IBP", "trig-sub" is beside the point right?

OpenStudy (amistre64):

with partial decomp; i get to: 3 ln(1-t^2) : but id have to derive it to see if i get back down :)

OpenStudy (amistre64):

droped a negative along the way i think

OpenStudy (anonymous):

lol, We just got lectured on IBP, so I think its more of a U sub problem as of right now.. What we did looks like a U sub, a very interesting one..

OpenStudy (anonymous):

The answer is Pi.

OpenStudy (amistre64):

\[\int \frac{3}{1-t}dt\ +\int\frac{3}{1+t}dt\] \[-3\int \frac{-1}{1-t}dt\ +3\int\frac{1}{1+t}dt\] \[-3ln(1-t) +3ln(1+t)\] \[3ln(1+t)-3ln(1-t)\] \[3[ln(1+t)-ln(1-t)]\] \[3\ ln\frac{(1+t)}{(1-t)}\]

OpenStudy (anonymous):

Hmm I put the definite integral in there and didn't get pi.. interesting.

OpenStudy (anonymous):

Okay Im going to ask the teacher about this one.. but anyway, Another quick question, I have another integral, the denominator is 1+16x^2 -- can I divide every part of the denom by 16 so I get 1/16 +x^2 if i put a 16 outside of the integral?

OpenStudy (amistre64):

as long as your modifying your integral by a useful form of "1", your fine

OpenStudy (amistre64):

or the lesser know, useful form of "0"

OpenStudy (anonymous):

Thanks